1. At a certain conference of 100 people,there are 29 Indians women and 23 Indian men. Of these Indian people 4 are doctors and 24 are either men or doctors. There are no foreign doctors. If 8A foreigners and B women doctors are attending the conference, then find the value of A-B.

Grade:10

1 Answers

Ravi
9 years ago
Number of indians= indian men+indian women
= 29+23=52

Number of foreigners=100-52=48

Given, number of foreigners=8A

So, A=6
And of the 24 people whose information is given as:
‘Of these Indian people 4 are doctors and 24 are either men or doctors.’

As we have 23 indian men, there are 3 who are men and doctors as:
The following eqn is just written considering only the Indian doctors and men specifically as information given in the question
n(M+D)= n(M)+n(D)-n(M.D)
24= 23+4-n(M.D)
n(M.D)=3

this means that 3 indian men who are doctors.

As 4 indians are doctors, so only 1 indian woman is a doctor

And , there is no foreign doctor, so
Number of women doctors at the conference=1 =B

B=1

So, A-B=6-1=5